A square wire loop with 3.00m sides is perpendicul
Answer #1

B(t) = (9.00 T/s) t - (3.60 T/sec)^{2} t^{2}

\phi (B) = B(t) . A

E = induced EMF = -d\phi _{B}/dt = - A dB(t)/dt

= - (1/2) X 3m X 3m X [ 9.00 - 7.2 t]

= -4,5 X [ 9-7.2t]

E_{t=2} = - 4.5 [9-7.2 X 2]= 24.3 Volts

Since dB/dt is negative , induced emf must oppose the reduction .

[Note:The diagram in this problem does not give exact direction of B field( It should normally be represented by X or . ) It is assumed B is out of the plane of the screen and since it is reducing, the induced current should be such as to produce B to increase in outward direction and hence should be anti-clockwise]

b . So total emf at t=2 is = 20v - 24.3 V = - 4.3V

c Total current will be anti-clockwise

d. Total current -4.3v/210 ohms = -0.20 Ameres